# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

### v2

@@ Introduction
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, following the final conditions of the solidification step. The initial condition for (31) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (31) without any significant error.
@@
Model for primary drying
In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [18]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated consecutively after the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
@@
where $T_{\textrm{c}}$ is the chamber wall temperature and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of (33). The mass balance of water at the sublimation front gives
@@
where $N_{\textrm{w}}$ is the sublimation flux and $\rho_{\textrm{e}}$ is the effective density of the dried region above the sublimation front. The driving force for mass transfer at the sublimation interface is [18, 1, 22]
@@
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation is described by [22]
@@
The variation of the mass transfer resistance can be approximated by the empirical expression [18, 4, 22]
@@
where $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, and $R_{\textrm{p2}}$ are the constants to be estimated from data. Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
The PDE (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
@@
Model for secondary drying
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3C), which is consistent with the primary drying model. The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated consecutively after the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
@@
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in (33), and the other parameters are as defined in (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
@@
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [15, 16, 4]
@@
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency), $E_{\textrm{a}}$ is the activation energy, and $R$ is the gas constant. The above desorption kinetics (43) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [15, 43, 19, 23, 4]. Further simplification that is relatively common can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in Ref. [16] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [4].
The governing PDE (42) requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying (38). Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big{(}T_{\textrm{u}}^{4}-T^{4}\big{)},\qquad z=0.$ (46)
The initial conditions for (42) and (43) are
@@
Modeling heat transfer in suspended vials
@@
Convection can be simply modeled using Newton’s law of cooling. Heat transfer coefficients for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref. [4].
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on Refs. [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; more detailed discussion can be found in the aforementioned references. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
@@
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, (50) can be simplified as
@@
In (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2). In this case, (51) becomes
@@
where $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among the vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, (52) needs to be modified to account for interactions between vials.
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref. [39]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref. [7]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify (52) to follow (49) as
@@
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (50). In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in (6), (32), and (42). If the chamber design and vial configuration are identical for both freezing and drying steps, the same value of $\mathcal{F}_{\textrm{s3}}$ can be used for all equations. A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in (39) and (46).
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of Ref. [31]), which is tiny. The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., (8), (38), and (45). This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
@@
Default model parameters
This section defines the default parameter values for the model developed in Sections 3.2, 3.3, and 3.4. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:Tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
@@
$\Delta H_{\textrm{des}}$ 2.68$\times$106 J/kg [16]
$\Delta H_{\textrm{fus}}$ 3.34$\times$105 J/kg [41]
$\Delta H_{\textrm{sub}}$ 2.84$\times$106 J/kg [19]
$\Delta H_{\textrm{vap}}$ see (14) J/kg [41]
@@
$\rho_{\textrm{e}}$ 215 kg/m3 [16]
$\rho_{\textrm{d}}$ 212.21 kg/m3 [16]
@@
$\rho_{\textrm{i}}$ 917 kg/m3 [32]
$\rho_{\textrm{s}}$ 1587.9 kg/m3 [32]
$\rho_{\textrm{w}}$ 1000 kg/m3 [13]
@@
$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [7]
$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [1]
@@
$p_{\textrm{w,c}}$ 3 Pa [7]
$p_{\textrm{w,sat}}$ see (12) and (36) Pa [41, 22]
@@
$H$ 7.2$\times$10-3 m calculated (see Appendix B)
$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix B)
$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix B)
$V_{\textrm{l}}$ 3$\times$10-6 m3 [7]
$x_{\textrm{s}}$ 0.05 [7]
@@
$\epsilon_{\textrm{gl}}$ 0.8 [31]
@@
$M_{\textrm{in}}$ 0.028 kg/mol [44]
$M_{\textrm{s}}$ 0.3423 kg/mol [13]
$M_{\textrm{w}}$ 0.018 kg/mol [44]
@@
Implementation
The mechanistic model for continuous lyophilization developed in this article is aimed to be used for various purposes, ranging process design and development to the actual operation (e.g., real-time control, state estimation, and optimization). In continuous operation, the process runs continuously with minimal human intervention, and thus efficient (ideally real-time) simulation and computation are important. Therefore, appropriate numerical methods must be applied. The detailed numerical methods and implementation are given in Appendix C. All simulations, calculations, and results presented in this article were performed and generated using MATLAB R2023a.
The model, data, and MATLAB code used in this work are made available (see the Data Availability section). In addition to the MATLAB implementation, we also provide an equivalent software package in Julia, a high-level, open-source programming language for scientific computing, for the selected simulations.
Results and Discussion
Model validation

### v3

@@ 3.2.5 Cooling
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in Equations (7), (8), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, following the final conditions of the solidification step. The initial condition for Equation (31) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from Equation (31) without any significant error.
@@
3.3 Model for primary drying
In conventional lyophilization, vials are placed on the heating shelf, and thus heat transfer at the bottom surface is driven by conduction (at the point of contact), convection (gas), and radiation, whereas only radiation dominates at the side and top surfaces [Pikal2005Model]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Figure 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated consecutively after the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
@@ 3.3 Model for primary drying
where $T_{\textrm{c}}$ is the chamber wall temperature and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of Equation (33). The mass balance of water at the sublimation front gives
@@
where $N_{\textrm{w}}$ is the sublimation flux and $\rho_{\textrm{e}}$ is the effective density of the dried region above the sublimation front. The driving force for mass transfer at the sublimation interface is [Pikal2005Model, Fissore2018Review, Bano2020LumpedDrying]
@@
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation is described by [Bano2020LumpedDrying]
@@
The variation of the mass transfer resistance can be approximated by the empirical expression [Pikal2005Model, Fissore2015Review, Bano2020LumpedDrying]
@@
where $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, and $R_{\textrm{p2}}$ are the constants to be estimated from data. Physically, $R_{\textrm{p0}}$ represents the resistance associated with mass convection above the product, which could include the presence of a skin layer on the cake surface and the effect of the stopper [Stratta2024Stopper], whereas $\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S}$ corresponds to the resistance associated with diffusion through the porous dried layer in the product.
Modeling the heat transfer and sublimation front in one dimension, as described in Equations (32) and (34), implies an assumption that the sublimation front is flat. This assumption has been widely used for conventional lyophilization as the heat flux through the bottom of the vial is much stronger than that from the sidewall, mainly due to heat conduction. In lyophilization of suspended vials, due to the absence of conductive heat transfer at the bottom surface, the sidewall heat flux becomes more comparable to the bottom heat flux. However, the bottom heat flux still dominates for two key reasons. First, the bottom shelf remains the primary heat source in the system, which directly interacts with the bottom surface of the vial through thermal radiation and convection. Second, since most vials are surrounded by others at similar temperatures, the driving force for sidewall heat transfer is significantly weaker than that for bottom heat transfer. A systematic analysis for comparing different heat transfer modes can be done using relevant dimensionless numbers, e.g., in Appendix A.
The PDE represented by Equation (32) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (38)
@@
3.4 Model for secondary drying
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Figure 3C), which is consistent with the primary drying model. The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated consecutively after the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
@@ 3.4 Model for secondary drying
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the concentration of bound water (aka moisture content, residual moisture, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the heat of desorption, $Q_{\textrm{rad}}$ is as defined in Equation (33), and the other parameters are as defined in Equation (32), with the subscript ‘e’ denoting the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
@@
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{d}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [Liapis1994Original, Sadikoglu1997Modeling, Fissore2015Review]
@@
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency), $E_{\textrm{a}}$ is the activation energy, and $R$ is the gas constant. The above desorption kinetics, Equation (43), is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [Liapis1994Original, Sircar2000LDF, Veraldi2008SimBatchModels, Fissore2011SecDryingMonitor, Fissore2015Review]. Further simplification that is relatively common can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in [Sadikoglu1997Modeling] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [Fissore2015Review].
The governing PDE, Equation (42), requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
$k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad z=H,$ (45)
where the value of $h_{\textrm{b}}$ is approximated to have the same value as that used in primary drying, Equation (38). Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\mathcal{F}_{\textrm{s1}}\big(T_{\textrm{u}}^{4}-T^{4}\big),\qquad z=0.$ (46)
The initial conditions for Equations (42) and (43) are
@@
3.5 Modeling heat transfer in suspended vials
@@
Convection can be simply modeled using Newton’s law of cooling. Heat transfer coefficients for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in [Fissore2015Review].
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on in [Mills1995HeatTransfer] and [Srisuma2024Rad]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; more detailed discussion can be found in the aforementioned references. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
@@
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, Equation (50) can be simplified as
@@
In Equation (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2). In this case, Equation (51) becomes
@@
where $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by Equation (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among the vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, Equation (52) needs to be modified to account for interactions between vials.
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in [Srisuma2024Rad]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by [Capozzi2019ContLyo_SuspendedVials]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify Equation (52) to follow Equation (49) as
@@
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using Equation (50). In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in Equations (6), (32), and (42). If the chamber design and vial configuration are identical for both freezing and drying steps, the same value of $\mathcal{F}_{\textrm{s3}}$ can be used for all equations. A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in Equations (39) and (46).
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of [Mills1995HeatTransfer]), which is tiny. The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., Equations (8), (38), and (45). This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., Equation (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
@@
3.6 Default model parameters
This section defines the default parameter values for the model developed in Sections 3.2, 3.3, and 3.4. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:ch2-tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
@@ 3.6 Default model parameters
[…]
$\Delta H_{\textrm{vap}}$ see Equation (14) J/kg [Smith2018Thermo]
@@
$\rho_{\textrm{e}}$ 215 kg/m3 [Sadikoglu1997Modeling]
$\rho_{\textrm{d}}$ 212.21 kg/m3 [Sadikoglu1997Modeling]
@@
$\rho_{\textrm{i}}$ 917 kg/m3 [Colucci2020CrystalSize]
$\rho_{\textrm{s}}$ 1587.9 kg/m3 [Colucci2020CrystalSize]
$\rho_{\textrm{w}}$ 1000 kg/m3 [Deck2024Freezing2D]
@@
$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [Capozzi2019ContLyo_SuspendedVials]
$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [Fissore2018Review]
@@
$p_{\textrm{w,c}}$ 3 Pa [Capozzi2019ContLyo_SuspendedVials]
$p_{\textrm{w,sat}}$ see Equations (12) and (36) Pa [Smith2018Thermo, Bano2020LumpedDrying]
@@
[…]
$x_{\textrm{s}}$ 0.05 [Capozzi2019ContLyo_SuspendedVials]
@@
$\epsilon_{\textrm{gl}}$ 0.8 [Mills1995HeatTransfer]
@@
$M_{\textrm{in}}$ 0.028 kg/mol [Chang2010Chem]
$M_{\textrm{s}}$ 0.3423 kg/mol [Deck2024Freezing2D]
$M_{\textrm{w}}$ 0.018 kg/mol [Chang2010Chem]
@@
4 Numerical Methods
This section comprehensively describes the numerical techniques required for solving the model equations efficiently. For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs). In [Deck2022FreezingLumped, Deck2024Freezing2D], the ODEs were solved numerically using an explicit scheme with fixed time steps. The probability of nucleation (Equation (24)) was calculated at every time step in a discrete fashion. From the computational perspective, this strategy is not the most efficient approach as it requires small step size to ensure stability. In this work, we rely on an implicit method with adaptive time steps; MATLAB’s ode15s is selected. Nevertheless, since Equation (24) is an algebraic equation that depends on the time step $\Delta t$, solving an ODE system coupled with Equation (24) requires fixing $\Delta t$ via a predefined time span and then running the solver repeatedly and sequentially for each time step, with the probability of nucleation calculated at the end of each run, until the first nucleation is detected. A more efficient strategy is to consider a continuous version of Equation (24), which is
$P=1-\exp\left(\int_{t_{0}}^{t}-\lambda dt\right),$ (54)
where $t_{0}$ is the initial time and $P(t=t_{0})=0$. Differentiate Equation (54) results in
$\frac{dP}{dt}=\lambda(1-P).$ (55)
The resulting ODE, Equation (55), can now be coupled with a system of ODEs describing heat and mass transfer in the process (e.g., during preconditioning and VISF), which can then be integrated using the selected ODE solver. In MATLAB, the Event function can be used to detect the time of first nucleation, i.e., when the probability exceeds the given value. With this strategy, our freezing model can be integrated continuously in a single run (i.e., without running the solver repeatedly). Consequently, the model equations can be solved within 0.1 s on a normal laptop, which is highly efficient.
For the PDEs in the primary drying and secondary drying models, the method of lines [Schiesser1991MOL] is recommended. The method of lines consists of two steps. First, the PDEs are discretized spatially to produce a system of ODEs. Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work). For the primary drying model, first consider Equation (32). As the equation involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (56)
where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, Equations (32), (38) and (39) become
$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (57)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (58)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (59)
To spatially discretize the PDE, define
$\Delta\xi=\frac{1}{n_{z}-1},$ (60)
$\xi=(j-1)\Delta\xi,$ (61)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing Equations (57)–(59) using the second-order finite difference scheme results in
$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (62)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (63)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (64)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in Equation (63) can be eliminated by substituting $T_{j=n_{z}+1}$ into Equation (62) for $j=n_{z}$. Similarly, $T_{j=0}$ in Equation (64) can be eliminated by substituting $T_{j=0}$ into Equation (62) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
$\Delta z=\frac{H}{n_{z}-1},$ (65)
$z=(j-1)\Delta z,$ (66)
where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$. Discretizing Equations (42), (43), (45), and (46) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (67)
$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (68)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (69)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (70)
All simulations, calculations, and results presented in this article were performed and generated using MATLAB R2023a. With the aforementioned numerical methods, our model can be simulated accurately within 1 s on a normal laptop. The model, data, and MATLAB code used in this work are made available (see the Data Availability section).
5 Results and Discussion
5.1 Model validation